The electric eel (Electrophorus electricus) can generate a potential difference of 600 V between a region just behind its head and its tail. Suppose an unsuspecting fish swims into this area, resulting in a lethal current of 1.1 A passing through the length of its body. If the fish is approximately 31 cm in length and 7.5 cm in diameter, what is its average resistivity?
Answer #1

The resistance within fish is

\(R=\frac{V}{I}\)

\(R=\frac{600}{1.1}\)

\(R=545.45 \Omega\)

now resistance formula in terms of length and area

\(R=\frac{\rho L}{A}\)

\(R=\frac{4 \rho L}{\pi d^{2}}\)

Resistivity of eel is

\(\rho=\frac{R \pi d^{2}}{4 L}\)

\(\rho=\frac{(545.45) \pi(0.075)^{2}}{4(0.31)}\)

\(\rho=7.773 \Omega m\)